# Bayesian inference and cure rate modeling for event history data

### v1

@@ 3.3 Metropolis-coupled MCMC algorithm
[…]
@@ 4 Controlling the False Discovery Rate
Let $D_{0}$ denote the total number of censored items in our sample, i.e., the cardinality of set $\Delta_{0}$. For ease of notation, we will rearrange the observed times so that the first $D_{0}$ subjects correspond to censored ones, that is, $\delta_{i}=0$ for $i=1,\ldots,D_{0}$, and $\delta_{i}=1$ for $i>D_{0}$. Within our Bayesian approach (see also 61, 62), it follows that
@@
Let now $q_{1}\geqslant\ldots\geqslant q_{D_{0}}$ denote the ordered values of $\widehat{\mathrm{P}}(I_{j}=0|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})$, for $j=1,\ldots,D_{0}$. Define $G_{j}:=\sum_{i=1}^{j}(1-q_{j})/j$, $j=1,\ldots,D_{0}$. For any given tolerance level $10<\alpha<$ 1, consider the decision rule
@@
where $k_{\alpha}=\max\{j=1,\ldots,D_{0}:G_{j}\leqslant\alpha\}$. It readily follows that $\hat{\mathbb{E}}(\mathrm{FDR}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})\leqslant\alpha$ (see 48, Section 2.6).
@@ 5 Results
Section 5.1 uses a large scale simulation study in order to evaluate the accuracy of the proposed method in terms of point estimation, assess the robustness to prior assumptions and also benchmark our results against the EM algorithm. The ability of controlling the FDR is also discussed. Finally, a real dataset application is presented in Section 5.2. Details of hyperparmeters of the prior distribution and the MCMC sampler are given in Appendix A. Further simulation results and insights are given in Appendix D.
@@ 5.1 Simulation study
Table 1: The parameter values, the deduced cure rate and censoring proportion (among susceptibles), for the simulated datasets.
@@
[…]
Next, we summarize our findings across all datasets and simulation scenarios. Our method is compared against the EM algorithm (the EM algorithm was used in [44]), based on a small-EM initialization procedure [63] with 45 random starts (the starting values of the EM algorithm were obtained under the same random initialization scheme as in the MCMC sampler; see Appendix A). Figure 4 compares point estimation accuracy for each parameter in terms of the Mean Absolute Error (MAE) between the true value and the point estimate according to each method. The bars represent the cumulative MAE (note that the contribution of each parameter is displayed using distinct colors). It is clear that our MCMC implementation is quite robust for both prior assumptions across all simulation scenarios. More specifically, the MAEs are almost equal for larger sample sizes ($n=2000$ or $n=5000$). In the case of smaller sample size ($n=500$), the regularized prior distribution tends to produce smaller MAEs when compared to the vague prior. Both MCMC implementations outperform the EM algorithm, despite the fact that we have used a rather large number (45) of small runs to initialize the final algorithm. Further summary results per scenario are shown in Figures 11, 12, 13, 14, 15, 16 and 17 in Appendix D.2. In almost all cases, the estimated marginal posterior distributions are multimodal and this mainly affects $\gamma$, $\beta_{0}$, $\beta_{1}$ and $\beta_{2}$. However, it is evident that the MAP estimate of each parameter is close to the “true” value.
@@ 5.2 Illustrative example: recidivism
Before implementing our method we have transformed age so the sample mean and variance equal 0 and 1, respectively. Our results are based on 4 separate runs of Algorithm 2 with different random starting values. For each run we considered $C=16$ tempered chains and a total of $N=100,000$ MCMC cycles. The temperature of each chain was selected according to the following scheme:
$h_{c}=\frac{1}{(1+\varepsilon)^{c^{d}-1}},\quad c=1,\ldots,C,$
where $\varepsilon=0.001$ and $d=2.2$. Our reported results are based on discarding the first $30,000$ cycles as burn-in period, and then thinning the chains by keeping every 10th sampled value.
@@ 6 Concluding remarks
A natural extension of the current study may be to explore issues regarding, for example, variable selection; within the Bayesian paradigm, this could be achieved by applying stochastic search variable selection techniques [65, 66] or incorporating shrinkage prior distributions [67]. Model adequacy may be another issue that should be checked; this could be addressed under a predictive viewpoint using cross validation [68], posterior predictive $p$-values [69, 70] or Bayesian goodness of fit-testing procedures [71]. Among authors’ future research directions is also the development of an R package, which would implement the proposed methodology.
@@ Appendix A Hyperparameters and MCMC sampler details
In order to select values of the scale parameters of our Metropolis-Hastings and MALA proposals with a reasonable acceptance rate betweeen proposed moves, the MCMC sampler runs for an initial warm-up period at Step 0 of Algorithm 2. During this period, each parameter is adaptively tuned as the MCMC sampler progresses in order to achieve acceptance rates of the proposed updates between pre-specified limits. The final value of each parameter is then selected as the one that will be used in the subsequent main MCMC sampler. More specifically, the scale parameter of the proposal distributions in Equations (15), (18), (20), (16), (22) were tuned in order to achieve acceptance rates of the proposed moves between $15\%-30\%$. The scale parameter ($\tau$) of the MALA proposal in Equation (10) was tuned until an acceptance rate between $40\%-60\%$ is achieved (see also 55).
@@
where $\varepsilon>0$, and $d>0$. In our simulations the values $\varepsilon=0.02$ and $b=2.3$ were used. The MCMC sampler was run for a total of $N=20000$ MCMC cycles, with each cycle consisting of $m_{1}=10$ MCMC iterations.
@@ B.1 Metropolis-Hastings step
$\tilde{\gamma}\sim\mathcal{N}\left(\gamma,s^{2}_{\gamma}\right).$ (14)
@@
Because the density of the proposal distribution in (14) is symmetric to $\gamma^{(t-1)}$ and $\tilde{\gamma}$, the Metropolis-Hastings acceptance probability reduces to
$A_{1,h}\left(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\gamma})}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\gamma)}\right\}.$ (15)
@@
$\tilde{\lambda}\sim\mathcal{LN}\left(\log\left(\lambda\right),s^{2}_{\lambda}\right).$ (16)
@@
The ratio of the densities of the proposal distribution in (16) is equal to
@@
$A_{2,h}\left(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\lambda})\tilde{\lambda}}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\lambda)\lambda}\right\}.$ (17)
@@
$\tilde{\alpha}_{1}\sim\mathcal{LN}\left(\log\left(\alpha_{1}\right),s^{2}_{\alpha_{1}}\right).$ (18)
@@
$A_{3,h}\left(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\alpha}_{1})\tilde{\alpha}_{1}}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\alpha_{1})\alpha_{1}}\right\}.$ (19)
@@
$\tilde{\alpha}_{2}\sim\mathcal{LN}\left(\log\left(\alpha_{2}\right),s^{2}_{\alpha_{2}}\right).$ (20)
@@
$A_{4,h}\left(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\alpha}_{2})\tilde{\alpha}_{2}}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\alpha_{2})\alpha_{2}}\right\}.$ (21)
@@
$\tilde{\boldsymbol{\beta}}\sim\mathcal{N}_{p}\left(\boldsymbol{\beta},\boldsymbol{\nu}\mathcal{I}_{p}\right)$ (22)
where $\boldsymbol{\nu}=(\tau_{0},\nu_{1},\ldots,\nu_{k})$ denotes a vector of fixed positive numbers corresponding to the variance of the proposal distribution. Because the density of the proposal distribution in (14) is symmetric to $\boldsymbol{\beta}$ and $\tilde{\boldsymbol{\beta}}$, the Metropolis-Hastings acceptance probability reduces to
$A_{5,h}\left(\boldsymbol{\beta},\tilde{\boldsymbol{\beta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\boldsymbol{\beta}})}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\boldsymbol{\beta})}\right\}.$ (23)

### v2

@@ 3.3 Metropolis-coupled MCMC algorithm
Algorithm 2 MALA-within-Gibbs MC3 targeting $\pi(\boldsymbol{\theta},\boldsymbol{I}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})$
@@ 4 Controlling the False Discovery Rate
Let $D_{0}$ denote the total number of censored items in our sample, i.e., the cardinality of set $\Delta_{0}$. For ease of notation, we will rearrange the observed times so that the first $D_{0}$ subjects correspond to censored ones, that is, $\delta_{i}=0$ for $i=1,\ldots,D_{0}$, and $\delta_{i}=1$ for $i>D_{0}$. Within our Bayesian approach (see also 55), it follows that
@@
Let now $q_{1}\geqslant\ldots\geqslant q_{D_{0}}$ denote the ordered values of $\widehat{\mathrm{P}}(I_{j}=0|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})$, for $j=1,\ldots,D_{0}$. Define $G_{j}:=\sum_{i=1}^{j}(1-q_{i})/j$, $j=1,\ldots,D_{0}$. For any given tolerance level $10<\alpha<$ 1, consider the decision rule
@@
where $k_{\alpha}=\max\{j=1,\ldots,D_{0}:G_{j}\leqslant\alpha\}$. It readily follows that $\hat{\mathbb{E}}(\mathrm{FDR}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})\leqslant\alpha$ (see 42, Section 2.6).
@@ 5 Results
Section 5.1 uses a large scale simulation study in order to evaluate the accuracy of the proposed method in terms of point estimation, assess the robustness to prior assumptions and also benchmark our results against the EM algorithm. The ability of controlling the FDR is also discussed. Finally, a real dataset application is presented in Section 5.2. Details of hyperparmeters of the prior distribution and the MCMC sampler are given in Section A of supplementary material. Further simulation results and insights are given in Section D of supplementary material.
@@ 5.1 Simulation study
Table 1: The parameter values, the deduced cure rate and censoring proportion (among susceptibles), for the simulated datasets. For each scenario and sample size level $n\in\{500,2000,5000\}$, a total of $N_{\mathrm{sim}}=20$ synthetic datasets were generated.
@@
At first we focus on the results for one typical simulated dataset of $n=500$ observations generated from Scenario A1 (see Table 1). Figure 2 shows the resulting estimates of the marginal posterior distribution for 4 different starting values of Algorithm 2, with $C=16$ tempered chains. The multi-modality in the marginal posterior distributions of $\gamma$, $\beta_{0}$, $\beta_{1}$ and $\beta_{2}$ is apparent. For this reason we will be using the Maximum A Posteriori (MAP) estimate as the point estimate of each parameter, i.e., the sampled parameter values $\boldsymbol{\theta}=(\gamma,\lambda,\alpha_{1},\alpha_{2},\beta_{0},\beta_{1},\beta_{2})$ which correspond to the iteration with the highest (joint) posterior density
$\hat{\boldsymbol{\theta}}=\mathrm{argmax}\{\pi(\boldsymbol{\theta}^{(t)}|\boldsymbol{y},\boldsymbol{\Delta},\boldsymbol{x}),t=1,\ldots,N\}.$
These correspond to the coloured vertical lines in Figure 2. We can see that the main mode of the posterior distribution (which corresponds to the one on top of the MAP estimate) contains the true parameter value. For interval estimation we will be using Highest (posterior) Density Intervals (HDI) (which could be discontinuous intervals in general). Observe also that all four runs of our MC3 sampler essentially lead to the same results, thus we can conclude that Algorithm 2 can successfully deal with this multimodal posterior distribution and produce a MCMC sample that converges in relatively small number of iterations in the posterior distribution. This is not the case for Algorithm 1, as shown in Figure 3, where the values of the observed log-likelihood are displayed for four randomly initialized chains. Observe that 2 among 4 chains (in particular chain 1 and chain 2) generated by Algorithm 1 are trapped in a minor mode (see also Section D.1 of supplementary material).
Next, we summarize our findings across all $N_{\mathrm{sim}}=20$ synthetic datasets and simulation scenarios. Our method is compared against the EM algorithm (the EM algorithm was used in [38]), based on a small-EM initialization procedure [56] with 45 random starts (the starting values of the EM algorithm were obtained under the same random initialization scheme as in the MCMC sampler; see Section A of supplementary material). Figure 4 compares point estimation accuracy for each parameter in terms of the Mean Absolute Error (MAE) between the true value and the point estimate according to each method. The MAE between the true parameter value $\theta_{j}\in\boldsymbol{\theta}=(\gamma,\lambda,\alpha_{1},\alpha_{2},\beta_{0},\beta_{1},\beta_{2})$ and the corresponding estimate for each specific synthetic dataset $\hat{\theta}_{j}$ is estimated as
$\mathrm{MAE}(\theta_{j},\hat{\theta}_{j})=\frac{1}{N_{\mathrm{sim}}}\sum_{i=1}^{N_{\mathrm{sim}}}|\theta_{j}-\hat{\theta}_{j}^{(i)}|,\quad j=1,\ldots,7$
where $\hat{\theta}_{j}^{(i)}$ denotes the estimate of $\theta_{j}$ arising from synthetic dataset $i$, $i=1,\ldots,N_{\mathrm{sim}}$.
The bars represent the cumulative $\mathrm{MAE}=\sum_{j=1}^{7}\mathrm{MAE}(\theta_{j},\hat{\theta}_{j})$ (note that the contribution of each parameter is displayed using distinct colors). It is clear that our MCMC implementation is quite robust for both prior assumptions across all simulation scenarios. More specifically, the MAEs are almost equal for larger sample sizes ($n=2000$ or $n=5000$). In the case of smaller sample size ($n=500$), the regularized prior distribution tends to produce smaller MAEs when compared to the vague prior. Both MCMC implementations outperform the EM algorithm, despite the fact that we have used a rather large number (45) of small runs to initialize the final algorithm. Further summary results per scenario are shown in Figures D.12, D.13, D.14, D.15, D.16, D.17 and D.18 in Section D.2 of supplementary material. In almost all cases, the estimated marginal posterior distributions are multimodal and this mainly affects $\gamma$, $\beta_{0}$, $\beta_{1}$ and $\beta_{2}$. However, it is evident that the MAP estimate of each parameter is close to the “true” value.
@@ 5.2 Illustrative example: recidivism
Before implementing our method we have transformed age so the sample mean and variance equal 0 and 1, respectively. Our results are based on 4 separate runs of Algorithm 2 with different random starting values. For each run we considered $C=16$ tempered chains and a total of $N=100,000$ MCMC cycles. Our reported results are based on discarding the first $30,000$ cycles as burn-in period, and then thinning the chains by keeping every 10th sampled value.
@@ 6 Concluding remarks
Our model is based on augmenting the data $\boldsymbol{y}=(y_{1},\ldots,y_{n})$ with the discrete latent cured status $\boldsymbol{I}=(I_{1},\ldots,I_{n})$ for each subject. Therefore, Hamiltonian MCMC methods as the No-U-Turn Sampler [58] implemented in the popular STAN software [59] are not directly applicable in this setting: only continuous parameters are allowed. Of course, one could integrate the discrete parameters ($\boldsymbol{I}$) out from the model and target the marginal posterior distribution (that is, $\pi(\boldsymbol{\theta}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})=\sum_{\boldsymbol{I}}\pi(\boldsymbol{\theta},\boldsymbol{I}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})$), but there are two drawbacks with this approach. At first, inference for the latent cured status would not be straightforward. Second, the multimodality of the posterior distribution would make the Hamiltonian Monte Carlo sampler getting trapped into minor modes and the proper exploration of the posterior surface would be questionable. In a similar manner, even if we integrate the discrete parameters out, approximate Bayesian techniques such as the Integrated Nested Laplace Approximation [60] are not expected to work well since the multimodal target posterior distribution would not justify to consider them at the first place.
It is also worth noting that, although we assume the time-to-event is subject to non-informative random right censoring, left-censored and/or interval-censored data can be seamlessly incorporated into our methodology. This is because the cured status would remain unknown only within the right-censored group, ensuring that crucial aspects of our approach stay consistent. It goes without saying that the detailed performance of the estimation algorithm and its numerical aspects, in the presence of such data, need to be evaluated. We plan to address this issue in future research.
Moreover, a natural extension of the current study may be to explore issues regarding, for example, variable selection; within the Bayesian paradigm, this could be achieved by applying stochastic search variable selection techniques [61, 62] or incorporating shrinkage prior distributions [63]. Model adequacy may be another issue that should be checked; this could be addressed under a predictive viewpoint using cross validation [64], posterior predictive $p$-values [65, 66] or Bayesian goodness of fit-testing procedures [67]. Among authors’ future research directions is also the development of an R package, which would implement the proposed methodology.
@@
  • We thank the Associate Editor and anonymous reviewers for their careful reading of our manuscript and their many insightful comments and suggestions.
Appendices
@@ Appendix A Hyperparameters and MCMC sampler details
In order to select values of the scale parameters of our Metropolis-Hastings and MALA proposals with a reasonable acceptance rate betweeen proposed moves, the MCMC sampler runs for an initial warm-up period at Step 0 of Algorithm 2. During this period, each parameter is adaptively tuned as the MCMC sampler progresses in order to achieve acceptance rates of the proposed updates between pre-specified limits. The final value of each parameter is then selected as the one that will be used in the subsequent main MCMC sampler. More specifically, the scale parameter of the proposal distributions were tuned in order to achieve acceptance rates of the proposed moves between $15\%-30\%$. The scale parameter ($\tau$) of the MALA proposal in Equation 11 was tuned until an acceptance rate between $40\%-60\%$ is achieved (see also 49).
@@
where $\varepsilon>0$, and $d>0$. In our simulations the values $\varepsilon=0.001$ and $d=2.5$ were used. The MCMC sampler was run for a total of $N=20000$ MCMC cycles, with each cycle consisting of $m_{1}=10$ MCMC iterations.
@@ B.1 Metropolis-Hastings step
$\tilde{\gamma}\sim\mathcal{N}\left(\gamma,s^{2}_{\gamma}\right).$ (B.15)
@@
Because the density of the proposal distribution in (B.15) is symmetric to $\gamma^{(t-1)}$ and $\tilde{\gamma}$, the Metropolis-Hastings acceptance probability reduces to
$A_{1,h}\left(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\gamma})}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\gamma)}\right\}.$ (B.16)
@@
$\tilde{\lambda}\sim\mathcal{LN}\left(\log\left(\lambda\right),s^{2}_{\lambda}\right).$ (B.17)
@@
The ratio of the densities of the proposal distribution in (B.17) is equal to
@@
$A_{2,h}\left(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\lambda})\tilde{\lambda}}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\lambda)\lambda}\right\}.$ (B.18)
@@
$\tilde{\alpha}_{1}\sim\mathcal{LN}\left(\log\left(\alpha_{1}\right),s^{2}_{\alpha_{1}}\right).$ (B.19)
@@
$A_{3,h}\left(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\alpha}_{1})\tilde{\alpha}_{1}}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\alpha_{1})\alpha_{1}}\right\}.$ (B.20)
@@
$\tilde{\alpha}_{2}\sim\mathcal{LN}\left(\log\left(\alpha_{2}\right),s^{2}_{\alpha_{2}}\right).$ (B.21)
@@
$A_{4,h}\left(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\alpha}_{2})\tilde{\alpha}_{2}}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\alpha_{2})\alpha_{2}}\right\}.$ (B.22)
@@
$\tilde{\boldsymbol{\beta}}\sim\mathcal{N}_{p}\left(\boldsymbol{\beta},\boldsymbol{\nu}\mathcal{I}_{p}\right)$ (B.23)
where $\boldsymbol{\nu}=(\tau_{0},\nu_{1},\ldots,\nu_{k})$ denotes a vector of fixed positive numbers corresponding to the variance of the proposal distribution. Because the density of the proposal distribution in (B.15) is symmetric to $\boldsymbol{\beta}$ and $\tilde{\boldsymbol{\beta}}$, the Metropolis-Hastings acceptance probability reduces to
$A_{5,h}\left(\boldsymbol{\beta},\tilde{\boldsymbol{\beta}}\right)=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\tilde{\boldsymbol{\beta}})}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{x})\dot{\pi}_{h}(\boldsymbol{\beta})}\right\}.$ (B.24)